Find the adjoint of F:C3 tends to C3 defined by F(z1, z2, z3)=(2z1+(1-i)z2 , (3+2i)z1 -4iz3 , 2iz1+(4-3i)z2 - 3z3 )


1
Expert's answer
2021-01-11T09:57:15-0500

Answer:


Step 1

F:43"\rarr"43 by

F"\lparen"z1,z2,z3"\rparen"= "\lparen"2 z1+"\lparen"1-i"\rparen"z2, "\lparen"3+2i"\rparen" z1-4iz3, 2iz1+ "\lparen"4-3i")"z2-3z3"\rparen"


Step 2

Way to find the adjoint of F

"("I) Find matrix of f with respect to basis "\lbrace"e1,e2, e3"\rbrace"

where e1="("1,0,0")" , e2="("0,1,0")" , e3="("0,0,1")"

Let A be the matrix corresponding to F.


 "("II) Compute "(" "\overline{ A^T}"")"


 "("II) Hence adjoint F*: 43"\rarr"43 is

 F* "\lparen"z1,z2,z3"\rparen"= "(" "\overline{ A^T}"")" "\begin{bmatrix}
z1\\
z2\\
z3\\
\end{bmatrix}"

Now

F"("e1")" = F"("1,0,0")" = "(" 2,"\lparen"3+2i"\rparen", 2i")"

F"("e2")" = F"("0,1,0")" = "(" 1-i, 0, 4-3i")"

F"("e3")" = F"("0,0,1")" = "(" 0,-4i, -3")"


Step 3

Hence


A= "\begin{bmatrix}
2 & 1-i & 0 \\
3+2i & 0 & -4i\\
2i & 4-3i & -3 \\
\end{bmatrix}"


AT= "\begin{bmatrix}
2 & 3+2i & 2i \\
1-i & 0 & 4-3i \\
0 & -4i & -3 \\
\end{bmatrix}"


 "\overline{ A^T}" ="\begin{bmatrix}
2 & 3-2i & -2i \\
1+i & 0 & 4+3i \\
0 & 4i & -3 \\
\end{bmatrix}"


Hence  F*: 43"\rarr"43 is

 F* "\lparen"z1,z2,z3"\rparen"= "(" "\overline{ A^T}"")" "\begin{bmatrix}
z1\\
z2\\
z3\\
\end{bmatrix}"

="\begin{bmatrix}
2 & 3-2i & -2i \\
1+i & 0 & 4+3i \\
0 & 4i & -3 \\
\end{bmatrix}""\begin{bmatrix}
z1\\
z2\\
z3\\
\end{bmatrix}"


="\lparen"2 z1+"\lparen"3+2i"\rparen"z2- 2iz3, "\lparen"1+i"\rparen"z1+4+3iz3, iz2-3z3"\rparen"

Therefore,

F "\lparen"z1,z2,z3"\rparen"="\lparen"2 z1+"\lparen"3+2i"\rparen"z2- 2iz3, "\lparen"1+i"\rparen"z1+4+3iz3, iz2-3z3"\rparen"